OpenStudy (anonymous):

Who can Read jsMath style and change it to"formattedText" type? For jsMath, it said "f(x) = \begin{cases} \ 2 & \text{ if } x \gt 1 \\ \ 4 & \text{ if } x =1 \\ \ -x+ 2 & \text{ if } -8 \le x \lt 1 \\ \ 10 & \text{ if } x \lt -8 \end{cases}" How to change it to "formattedText" type?

OpenStudy (anonymous):

f(x) = \begin{cases} \ 9 & \text{ if } x \gt 8 \\ \ -3 & \text{ if } x =8 \\ \ -x+ 9 & \text{ if } -4 \le x \lt 8 \\ \ 13 & \text{ if } x \lt -4 \end{cases}

OpenStudy (anonymous):

f(x) = \begin{cases} \ 9 & \text{ if } x \gt 8 \\ \ -3 & \text{ if } x =8 \\ \ -x+ 9 & \text{ if } -4 \le x \lt 8 \\ \ 13 & \text{ if } x \lt -4 \end{cases}